In an RLC series circuit that includes a source of alternating current operating at fixed frequency and voltage, the resistance R is equal to the inductive reactance. The RLC series circuit uses a parallel plate capacitor. If the plate separation of the capacitor is reduced to half of its original value, the current in the circuit doubles. Find the initial capacitive reactance in terms of R.
Answer #1

XL = R

wL = R     

lets assume capacitive reactance = a         

Zinitial = sqrt (R2 + (R- a)2 )                           eq-1

current in circuit doubles since impedence gets halved

so Zfinal = (0.5) Zinitial

sqrt (R2 + (R- a/2)2 ) = (0.5) sqrt (R2 + (R- a)2 )

(R2 + (R- (a/2))2 ) = 0.25 (R2 + (R- a)2 )

R2 + R2 + a2/4 - Ra = 0.25 (R2 + R2 + a2 -2 Ra )

2R2 + a2/4 - Ra = 0.5 R2 + 0.25 a2 - 0.5 R a

1.5 R2 = 0.5 R a

a = 3R